# SpecEval: Evaluating Code Comprehension in Large Language Models via Program Specifications

### v1

@@ Abstract.
Large Language models (i.e. LLMs) have achieved impressive performance in automated software engineering. Extensive efforts have been made to evaluate the abilities of code LLMs in various aspects, with an increasing number of benchmarks and evaluation frameworks proposed. Apart from the most sought-after capability of code generation, the capability of code comprehension is being granted growing attention. Nevertheless, existing works assessing the code comprehension capability of LLMs exhibit varied limitations. For example, evaluation frameworks like CRUXEval and REval usually focus on code reasoning tasks over a certain input case. It leads to a limited range of execution traces covered, resulting in a loss in code semantics examined and the inability to assess the comprehensive understanding of LLMs concerning the target program.
To tackle the aforementioned challenges, we propose SpecEval, a novel black-box evaluation framework to evaluate code comprehension in LLMs via program specifications. Inspired by the idea that specifications can act as a comprehensive articulation of program behaviors concerning all possible execution traces, we employ formalized program specifications to represent program semantics and perform comprehensive evaluations. In particular, four specification-related tasks are designed meticulously to assess the capability of LLMs from basic to advanced levels. Moreover, counterfactual analysis is conducted to study the performance variance of LLMs under semantics-preserving perturbations, and progressive consistency analysis is performed to study the performance consistency of LLMs over a series of tasks with sequential dependence. Systematic experiments are conducted on six state-of-the-art LLMs. Extensive experimental results present a below-satisfactory performance of LLMs on specification-related tasks, revealing the limitations of existing LLMs in terms of articulating program semantics. The counterfactual analysis and progress consistency analysis also reveal the sensitivity of LLMs towards semantic-preserving perturbations and limited progressive consistency across the progressive tasks. Our experimental results and analysis underscore future directions for enhancement.
Large language models, Program Semantics, Program Specifications
ccs: General and reference Empirical studiesccs: Theory of computation Program specifications
@@ 1. Introduction
Automated software engineering, encompassing areas such as code generation (Liu et al., 2024b; Chen et al., 2021), program repair (Xia et al., 2023; Nilizadeh et al., 2021), and code summarization (Zhang et al., 2020; McBurney and McMillan, 2015), has persistently remained a prominent research focus in both academic and industrial practice. Early research in this field adopted traditional techniques, such as template-based (Flanagan and Leino, 2001; Ernst et al., 2007) and retrieval-based (Hayati et al., 2018; Zhang et al., 2020) methods, to address these issues. However, due to technology and data availability constraints, these methods often struggled to deliver satisfactory performance. With the rise and advancement of deep learning technologies, an increasing number (Yu et al., 2019), Transformer (Vaswani, 2017), and GNNs (Wu et al., 2020), have been applied to this field. Extensive research on program semantics learning with advanced neural network models drives significant progress and advancements in various SE applications. Program semantics are the essence of a program, programs with identical semantics may have different expressions. Consequently, numerous studies (Susanto and Lu, 2017; Zhou et al., 2019; Peng et al., 2021) have explored different techniques, including tree-based and graph-based approaches, to improve the neural networks in learning program semantics.
@@
Recently, large language models (LLMs) (OpenAI, 2023; Achiam et al., 2023; Wei et al., 2023; Nijkamp et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Touvron et al., 2023; Roziere et al., 2023) have impressed the community with their outstanding performance, vastly surpassing previous techniques in various software development applications. In the meantime, a wide range of software development tools (Github, 2024; JeyBrains, 2024) powered by these models have significantly increased the productivity of software developers. Behind the rapid advancement of these LLMs, compared to earlier techniques such as graph-based approaches that explicitly incorporated code semantics into model training to help the model comprehend program semantics, LLMs usually treat the code snippet as pure text and rely on the straightforward pre-training techniques i.e., next-token prediction to train on a massive amount of code corpus. As a result, it raises a question i.e., does the simple next-token pertaining technique truly enable models to comprehend the complex semantics of code?
@@
Numerous benchmarks, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and SWE-bench (Jimenez et al., 2023), are proposed to evaluate the coding capacity of LLMs. Although these benchmarks can, to some extent, reflect the code generation capabilities of different models, using the criterion of whether the generated code passes test cases is still insufficient to determine whether the models genuinely understand the code (Yang et al., 2024). To conduct a more in-depth analysis, CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024) define certain program execution tasks, such as code coverage prediction and output prediction, to assess the code reasoning capabilities of LLMs by comparing the model’s generated results with actual program runtime outcomes. Nevertheless, the dynamic execution results of a program depend on variations in its input, making it challenging to ensure that a comprehensive set of inputs can cover all possible execution traces  [^1] The execution trace refers to the transitions of program states (Plotkin, 1981) where the states are defined as the set of all program variables including their types and values.. It leads to limited program inputs invoking limited execution traces for a program and cannot comprehensively reflect the full program semantics. Some probing techniques (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) are adopted to analyze the code understanding, such as code syntax and semantic knowledge. They typically require access to a model’s intermediate vector representations for corresponding inputs to investigate whether the model has learned specific attributes. As these techniques usually require direct access to the model, they are not feasible for black-box models such as GPT-3.5 and GPT-4.
@@
These challenges inspire us to explore other forms of representation for program semantics, which are required to articulate program behaviors in multiple granularities. Formalized program specifications encompass precise claims that describe the intended or actual program behaviors in formal languages. By specifying the constraints on program variables or the transition of program states, the strong program specifications can effectively articulate program semantics, accurately capturing program behaviors either generally or in detail. Moreover, the correctness of formal specifications can be effectively validated by corresponding specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015). It also overcomes the limitations of using natural languages, such as code summaries, to describe the functionality of code, which cannot be automatically verified.
@@
Inspired by these ideas, we propose SpecEval, a novel black-box evaluation framework featuring the representation of program semantics in the form of formal program specifications. We define multiple specification-related tasks for subject LLMs to complete. These tasks require LLMs to demonstrate an understanding of code by summarizing the code behavior and generating formalized statements. By leveraging automated procedures for determining the correctness of specifications, the assessment can be conducted rigorously and efficiently. In particular, four tasks with sequential dependencies are designed, including Judgement, Selection, Infilling, and Generation. In the Judgement task, LLMs are required to evaluate the correctness of the given specifications for the code. The Selection task involves selecting the most appropriate candidate specification for the target program. The Infilling task requires LLMs to complete partially provided specifications by filling in the correct sub-expressions. Finally, the Generation task challenges LLMs to generate correct specifications entirely from scratch. For each task, we further introduce counterfactual analysis to cast code comprehension as the problem of determining how controlled input code changes result in model output changes by the semantic-preserving perturbations, with mutation operators including Def-use Break, If-else Flip, Independent Swap, Name Random, and Name Shuffle. Moreover, for each kind of perturbation, Progressive Consistency Analysis is conducted with the well-designed metric Progressive Consistency Score to examine the consistency performance of LLMs over these tasks with progressive dependencies, enabling the framework to assess the code comprehension capacities of LLMs from the shallow to the deep.
To construct the framework, we collect 204 Java programs with verifiable ground-truth specifications in JML (Leavens et al., 1998) style from existing datasets, including Frama-C-Problems (github, 2024), the Java category of SV-COMP Benchmark (sosy lab, 2024), and a dataset collected by Ma et al.(Ma et al., 2024). Test cases for these programs are also generated to assist the specification validation process. We conduct experiments on six large language models, including state-of-the-art code LLMs and general LLMs, for a systematic comparison between the popular models. Experiment results show that LLMs still have a long way to go before they can fully articulate program semantics with formal specifications, with less than satisfactory performance in most specification-related tasks. Also, Counterfactual Analysis discloses the sensitivity of LLMs to different extents towards semantic-preserving perturbations. Additionally, Progressive Consistency Analysis suggests that the reasoning ability of LLMs does not work in a progressive manner as humans do.
@@
  • A novel black-box evaluation framework assessing the program semantics learning abilities of LLMs, featuring the representation of program semantics in the form of formalized specifications. Four tasks with progressive difficulties and Counterfactual analysis are designed to study the performance of LLMs from multiple aspects.
  • An adapted benchmark (Author, s) based on existing datasets (sosy lab, 2024; github, 2024; Nilizadeh et al., 2021; Ma et al., 2024), consisting of 204 Java programs, along with ground-truth specifications, test cases for dynamic validation, and five semantics-equivalent variants, is released to facilitate follow-up research.
@@ 2.1. Program Specification
Program specification refers to precise statements that define a program’s intended or actual behaviors, either in its entirety or in specific components. According to the language adopted, program specifications can be categorized into natural language specifications and formalized specifications. Natural language specifications primarily encompass software documentation and code comments. Another significant portion of program specifications is articulated using formal languages, such as mathematical expressions (Leavens et al., 1998; Kirchner et al., 2015) to specify the constraints governing a program’s behavior, and automata (Lo and Khoo, 2006; Shoham et al., 2007) to describe the transitions of program states. Compared with natural language, formalized specifications are rigorous and unambiguous. Program specifications have been widely adopted in a variety of software engineering tasks, such as requirement engineering (Aman and Ibrahim, 2013; Sharifi et al., 2020), software testing (Mesbah et al., 2012; Nguyen and Dang, 2023), and model checking (Cabodi et al., 2008; Beyer et al., 2015).
In SpecEval, we mainly focus on formal specifications written in Java Modeling Language (JML) (Leavens et al., 1998) describing the actual behaviors of Java programs. One example program isPalindrome and corresponding ground-truth specifications are illustrated in Fig. 1. The program aims to check whether a given input string s is a palindrome string, i.e., a string that reads the same forward and backward. The specifications are instrumented within the program in the form of comments starting with //@. Generally, the specifications involved can be divided into three categories:
  • Preconditions (line 1), described in requires statements, specifying the constraints on input variables that must be met before the corresponding code (i.e. isPalindrome here) is executed.
  • Postconditions (line 2), described in ensures statements, specifying the properties of the returned value that always hold after the corresponding code is executed.
  • Loop Invariants (lines 5, 6, and 7), described in maintaining statements, specifying the constraints on the local variables that always hold each time before the loop body is executed.
Together, the specifications mentioned above constitute a detailed articulation of the semantics of the target program. For the method isPalindrome as a whole, preconditions and postconditions specify its overall requirements and functionalities. For the internal implementation of isPalindrome, loop invariants reveal the pattern followed by local variables when their values are updated.
@@
It is worth noting that formalized program specifications are more rigorous, eliminating ambiguity compared with specifications in natural language. Moreover, formalized program specifications can thoroughly articulate program semantics. Program semantics essentially lies in the evolution of program states, which are usually defined by the set of variable values. By specifying the constraints on program variables or the transition of program states, the strong program specifications are rich in semantic information and can accurately capture program behaviors. In addition, the correctness of formal specifications can be effectively validated. A good number of specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015) is proposed to automatically check the consistency between the given specifications and the target program. Consequently, compared with natural language specifications, the correctness of program specifications can be automatically verified.
2.2. Program Semantics Comprehension of LLMs
Program semantics are the core of a program, and only when code models can effectively learn these semantics can they provide reliable support and assurance in software development. Recently, there has been a surge in the impact of LLMs, playing a significant role in various software development scenarios. In light of the critical role of code semantics, a substantial body of research has attempted to evaluate whether LLMs genuinely comprehend code through various analytical techniques.
A typical technique involves using probing analysis (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) to explore whether code models have learned specific code attributes from the trained code datasets. Specifically, it leverages the transformation of code semantic information, such as data flow, into high-dimensional vector representations in the latent space. This enables an examination of whether the model truly understands the semantics of the code. Yet, they are usually required to access the model’s intermediate vector representations for analysis, which is not feasible for black-box commercial models such as GPT-3.5 and GPT-4. Moreover, some popular code generation benchmarks such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and SWE-bench (Jimenez et al., 2023) are proposed to evaluate the coding capacity of LLMs. Yet, the criterion to judge whether the generated code can pass test cases is insufficient to articulate the model’s understanding of the code.
The latest research REval (Chen et al., 2024) proposed a framework to evaluate the code reasoning ability of LLMs by inferring the program runtime behaviors. Inspired by CRUXEval (Gu et al., 2024), REval proposed four evaluation tasks, including Code Coverage Prediction, Program State Prediction, Execution Path Prediction, and Output Prediction for evaluation. In particular, Code Coverage Prediction requires LLMs to predict whether a statement will be executed given a certain input. Program State Prediction requires LLMs to reason about the type and value of a specific variable. Execution Path Prediction requires LLMs to predict the next statement executed under certain circumstances, and Output Prediction requires LLMs to generate the program’s output directly from the given input.
@@ 2.2. Program Semantics Comprehension of LLMs
It should be emphasized that in each of the four tasks mentioned, LLMs are provided with one specific input respectively. For a deterministic program, one specific input can only result in one particular trace of program state transition, i.e. the execution trace. This observation leads to the fact that for each independent task assigned, LLMs are only reasoning about program behaviors within one specific execution trace, whereas there exists an infinite number of execution traces (corresponding to the infinite number of inputs) within the program state space other than the one being examined. Nevertheless, program semantics is not only about the program behaviors during one specific execution but also about the general pattern of program state transition adaptable to all potential program states. The set of program behaviors for all potential program states and execution traces constitutes full program semantics. This indicates that the practice employed by REval could cause a loss in the program semantics examined by its tasks, since REval can only supply a limited number of evaluation tasks and inputs, leaving the other program behaviors in an infinite number of execution traces uncovered.

### v2

@@ Abstract.
Large Language models (i.e., LLMs) have achieved impressive performance in automated software engineering. Extensive efforts have been made to evaluate the abilities of code LLMs in various aspects, with an increasing number of benchmarks and evaluation frameworks proposed. Apart from the most sought-after capability of code generation, the capability of code comprehension is being granted growing attention. Nevertheless, existing works assessing the code comprehension capability of LLMs exhibit varied limitations. Evaluation frameworks like CRUXEval and REval usually focus on code reasoning tasks over a certain input case, leading to a limited range of execution traces covered, resulting in a loss in code semantics examined and the inability to assess the comprehensive understanding of LLMs concerning the target program.
To tackle these challenges, we propose SpecEval, a novel black-box evaluation framework to evaluate code comprehension in LLMs via program specifications. Inspired by the idea that specifications can act as a comprehensive articulation of program behaviors concerning all possible execution traces, we employ formalized program specifications to represent program semantics and perform comprehensive evaluations. In particular, four specification-related tasks are designed meticulously to assess the capability of LLMs from basic to advanced levels. Counterfactual analysis is further conducted to study the performance variance of LLMs under semantics-preserving perturbations. Systematic experiments are conducted on six state-of-the-art LLMs. Extensive experimental results present a below-satisfactory performance of LLMs on specification-related tasks, revealing the limitations of existing LLMs in terms of articulating program semantics with formal specifications. Counterfactual analysis also reveals the sensitivity of LLMs towards semantic-preserving perturbations.
@@ 1. Introduction
Automated software engineering, encompassing areas such as code generation (Liu et al., 2024c; Chen et al., 2021), program repair (Xia et al., 2023; Nilizadeh et al., 2021; Kong et al., 2024), and code summarization (Liu et al., 2020; McBurney and McMillan, 2015), has persistently remained a prominent research focus in both academic and industrial practice. Early research in this field adopted traditional techniques, such as template-based (Flanagan and Leino, 2001; Ernst et al., 2007) and retrieval-based (Hayati et al., 2018; Liu et al., 2020) methods, to address these issues. However, due to technology and data availability constraints, these methods often struggled to deliver satisfactory performance. With the rise of deep learning, techniques including RNN (Yu et al., 2019), Transformer (Vaswani, 2017), and GNNs (Wu et al., 2020) have been applied to this field. Extensive research on program semantics learning with advanced neural network models drives significant advancements in various SE applications. Program semantics are the essence of a program, programs with identical semantics may have different expressions. Consequently, numerous studies (Susanto and Lu, 2017; Zhou et al., 2019; Peng et al., 2021) have explored different techniques, including tree-based and graph-based approaches, to improve the neural networks in learning program semantics.
@@
Large language models (LLMs) (OpenAI, 2023; Achiam et al., 2023; Wei et al., 2023; Nijkamp et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Guo et al., 2025) have recently impressed the community with their outstanding performance, vastly surpassing previous techniques in various software engineering tasks. Meanwhile, a wide range of software development tools (Github, 2024; JeyBrains, 2024) powered by these models have significantly increased the productivity of software developers. Compared to earlier techniques such as graph-based ones that explicitly incorporated code semantics into model training to help the model comprehend program semantics, LLMs usually treat the code snippet as pure text and rely on the straightforward pre-training technique, i.e., next-token prediction, to train on a massive amount of code corpus. This raises a question: does next-token pertaining technique truly enable models to comprehend the complex code semantics?
@@
Numerous benchmarks, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and SWE-bench (Jimenez et al., 2023), are proposed to evaluate the coding capacity of LLMs. Although these benchmarks can, to some extent, reflect the code generation capabilities of different models, using the criterion of whether the generated code passes test cases is still insufficient to determine whether the models genuinely understand the code (Yang et al., 2024b). For more in-depth analysis, CRUXEval (Gu et al., 2024) and REval (Chen et al., 2024a) define certain program execution tasks, such as code coverage prediction and output prediction, to assess the code reasoning capabilities of LLMs by comparing the model’s generated results with actual program runtime outcomes. Nevertheless, the dynamic execution results of a program depend on its input, making it challenging to ensure that a comprehensive set of inputs can cover all possible execution traces  [^1] The execution trace refers to the transitions of program states (Plotkin, 1981) where the states are defined as the set of all program variables including their types and values.. This leads to limited program inputs invoking limited execution traces for a program and cannot comprehensively reflect the full program semantics. Some probing techniques (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) are adopted to analyze the code understanding, such as code syntax and semantic knowledge. They typically require access to a model’s intermediate vector representations for corresponding inputs to investigate whether the model has learned specific attributes. As these techniques usually require direct access to the model, they are not feasible for black-box models such as GPT-3.5 and GPT-4.
@@
These challenges inspire us to explore other forms of representation for program semantics, which are required to articulate program behaviors in multiple granularities. Formalized program specifications encompass precise claims that describe the intended or actual program behaviors in formal languages. By specifying the constraints on program variables or the transition of program states, the strong program specifications can effectively articulate program semantics, accurately capturing program behaviors either generally or in detail. Moreover, the correctness of formal specifications can be automatically and effectively validated by corresponding specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015), overcoming the limitations of natural language specifications such as code summaries.
@@
Inspired by these ideas, we propose SpecEval, a novel black-box evaluation framework featuring the representation of program semantics in the form of formal program specifications. We define multiple specification-related tasks for subject LLMs to complete. These tasks require LLMs to demonstrate an understanding of code by summarizing code behavior into formalized statements. Leveraging automated procedures for determining the correctness of specifications, the assessment can be conducted rigorously and efficiently. Specifically, four specification-related tasks are designed, including Judgement, Selection, Infilling, and Generation. In the Judgement task, LLMs are required to judge the correctness of the given specification. The Selection task involves selecting the most appropriate candidate specification for the target program. The Infilling task requires LLMs to complete partially provided specifications by filling in the correct sub-expressions. Finally, the Generation task challenges LLMs to generate correct specifications entirely from scratch. For each task, we further introduce counterfactual analysis to cast code comprehension as the problem of determining how controlled input code changes result in model output changes by the semantic-preserving perturbations, with mutation operators including Def-use Break, If-else Flip, Independent Swap, Name Random, and Name Shuffle. For evaluation, we collect 204 Java programs with verifiable ground-truth specifications in JML (Leavens et al., 1998) style from existing datasets. Test cases for these programs are also generated to assist the specification validation process. We conduct experiments on six large language models, including state-of-the-art code LLMs and general LLMs, for a systematic comparison between the popular models. Experimental results show that LLMs still have a long way to go before being able to fully articulate program semantics with formal specifications, with less than satisfactory performance in most specification-related tasks. Furthermore, Counterfactual analysis discloses the sensitivity of LLMs to different extents towards semantic-preserving perturbations.
@@
  • A novel black-box evaluation framework assessing the program semantics learning abilities of LLMs, featuring the representation of program semantics in the form of formalized specifications. Four different tasks and Counterfactual analysis are designed to study the performance of LLMs from multiple aspects.
  • An adapted benchmark based on existing datasets (sosy lab, 2024; github, 2024; Nilizadeh et al., 2021; Ma et al., 2024), consisting of 204 Java programs along with five semantics-equivalent variants, as well as ground-truth specifications.
@@ 2.1. Program Specification
Program specification refers to precise statements that define a program’s intended or actual behaviors. According to the language adopted, program specifications can be categorized into natural language specifications and formalized specifications. Natural language specifications primarily encompass software documentation and code comments. Another significant portion of program specifications is articulated using formal languages, such as mathematical expressions (Leavens et al., 1998; Kirchner et al., 2015) to specify the constraints on program behaviors, and automata (Lo and Khoo, 2006; Shoham et al., 2007) to describe the transitions of program states. Formalized specifications have been widely adopted in automated software engineering tasks, such as requirement engineering (Aman and Ibrahim, 2013; Sharifi et al., 2020), software testing (Mesbah et al., 2012; Nguyen and Dang, 2023), and model checking (Cabodi et al., 2008; Beyer et al., 2015).
In SpecEval, we mainly focus on formal specifications written in Java Modeling Language (JML) (Leavens et al., 1998) describing the actual behaviors of Java programs. One example program isPalindrome and corresponding ground-truth specifications are illustrated in Fig. 1. The program aims to check whether a given input string s is a palindrome string, i.e., a string that reads the same forward and backward. The specifications are instrumented within the program in the form of comments starting with //@. Generally, the specifications involved can be divided into three categories: Preconditions (Line 1), Postconditions (Line 2), and Loop Invariants (Line 5, 6, and 7). Together, the specifications mentioned above constitute a detailed articulation of the semantics of the target program.
@@
It is worth noting that formalized program specifications are more rigorous, eliminating ambiguity compared with specifications in natural language. Moreover, formalized program specifications can thoroughly articulate program semantics. Program semantics essentially lies in the evolution of program states, which are usually defined by the set of variable values. By specifying the constraints on program variables or the transition of program states, strong program specifications are rich in semantic information and can accurately capture program behaviors. In addition, the correctness of formal specifications can be effectively validated by specification verifiers (Cok, 2011; Ahrendt et al., 2014; Kirchner et al., 2015), which can automatically check the consistency between the given specifications and the target program. Consequently, compared with natural language specifications, the correctness of program specifications can be automatically verified.
2.2. Motivation
Program semantics are the core of a program, and only when code models can effectively learn these semantics can they provide reliable support and assurance in software development. In light of the critical role of code semantics, a substantial body of research has attempted to evaluate whether LLMs genuinely comprehend code through various analytical techniques. A typical white-box technique involves probing analysis (Hernández López et al., 2022; Wan et al., 2022; Ma et al., ) to explore whether code models have learned specific code attributes from the trained code datasets. It leverages the transformation of code semantic information, such as data flow, into high-dimensional vector representations in the latent space, enabling an examination of whether the model truly understands the semantics of the code. Another typical black-box framework REval (Chen et al., 2024a) evaluates the code reasoning ability of LLMs by inferring the program runtime behaviors. Inspired by CRUXEval (Gu et al., 2024), REval designs four evaluation tasks: Code Coverage Prediction requires LLMs to predict whether a statement will be executed given a certain input, Program State Prediction requires LLMs to reason about the value of a specific variable, Execution Path Prediction requires LLMs to predict the next statement to be executed, and Output Prediction requires LLMs to generate the program’s output directly from the given input.
@@ 2.2. Motivation
However, existing works face various limitations. Probing analysis usually requires access to the model’s intermediate vector representations for analysis, which is not feasible for black-box commercial models such as GPT-3.5 and GPT-4. As for the black-box approach REval, it should be noted that in each of the four tasks involved, LLMs are provided with one specific input respectively. For a deterministic program, one specific input can only result in one particular trace of program state transition, i.e. the execution trace. This observation leads to the fact that for each independent task assigned, LLMs are only reasoning about program behaviors within one specific execution trace, whereas there exists an infinite number of execution traces (corresponding to the infinite number of inputs) within the program state space other than the one being examined. Nevertheless, program semantics is not only about the program behaviors during one specific execution but also about the general pattern of program state transition adaptable to all potential program states. The set of program behaviors for all potential program states and execution traces constitutes full program semantics. This indicates that the practice employed by REval could cause a loss in the program semantics examined, since the evaluation tasks and inputs are limited, leaving the other program behaviors in an infinite number of execution traces uncovered.